# Raghav Sahidullah

### 02 Mar 2026 v1

@@ Abstract
This report presents the TCG CREST system description for Track 1 (Speaker Diarization) of the DISPLACE-M challenge, focusing on naturalistic medical conversations in noisy rural-healthcare scenarios. Our study evaluates the impact of various voice activity detection (VAD) methods and advanced clustering algorithms on overall speaker diarization (SD) performance. We compare and analyze two SD frameworks: a modular pipeline utilizing SpeechBrain with ECAPA-TDNN embeddings, and a state-of-the-art (SOTA) hybrid end-to-end neural diarization system, Diarizen, built on top of a pre-trained WavLM. With these frameworks, we explore diverse clustering techniques, including agglomerative hierarchical clustering (AHC), and multiple novel variants of spectral clustering, such as SC-adapt, SC-PNA, and SC-MK. Experimental results demonstrate that the Diarizen system provides an approximate $39\%$ relative improvement in the diarization error rate (DER) on the post-evaluation analysis of Phase I compared to the SpeechBrain baseline. Our best-performing submitted system employing the Diarizen baseline with AHC employing a median filtering with a larger context window of $29$ achieved a DER of 10.37% on the development and 9.21% on the evaluation sets, respectively. Our team ranked sixth out of the 11 participating teams after the Phase I evaluation.
@@ 2.2 Diarizen System
The Diarizen  [^4] https://github.com/BUTSpeechFIT/Diarizen/tree/main [12] system is the state-of-the-art SD system built on end-to-end neural diarization with vector clustering (EEND-VC) [13, 14, 15]. It is a hybrid speaker diarization pipeline. At the micro-level, it utilizes a powerful local end-to-end neural module, composed of a pre-trained WavLM [16] feature extractor, a Conformer [17] block, and a linear classifier, to process raw audio in short, overlapping chunks of typically $80$ to $16$ second windows. This acts as a neural frontend, which generates frame-level speech probabilities and extracts highly discriminative speaker embeddings while handling overlapping speech.
@@ 5 Conclusions and Future Directions
Although our best system ranked sixth among 11 unique teams on the leaderboard  [^8] Codabench DISPLACE-M Challenge Results of the DISPLACE-M Challenge under the Diarizen Evaluation Phase 1, we identify several limitations in our study and outline multiple promising directions for future work.

### 07 Mar 2026 v3

@@ Abstract
This report presents the TCG CREST system description for Track 1 (Speaker Diarization) of the DISPLACE-M challenge, focusing on naturalistic medical conversations in noisy rural-healthcare scenarios. Our study evaluates the impact of various voice activity detection (VAD) methods and advanced clustering algorithms on overall speaker diarization (SD) performance. We compare and analyze two SD frameworks: a modular pipeline utilizing SpeechBrain with ECAPA-TDNN embeddings, and a state-of-the-art (SOTA) hybrid end-to-end neural diarization system, Diarizen, built on top of a pre-trained WavLM. With these frameworks, we explore diverse clustering techniques, including agglomerative hierarchical clustering (AHC), and multiple novel variants of spectral clustering, such as SC-adapt, SC-PNA, and SC-MK. Experimental results demonstrate that the Diarizen system provides an approximate $39\%$ relative improvement in the diarization error rate (DER) on the post-evaluation analysis of Phase I compared to the SpeechBrain baseline. Our best-performing submitted system employing the Diarizen baseline with AHC employing a median filtering with a larger context window of $29$ achieved a DER of 10.37% on the development and 9.21% on the evaluation sets, respectively. Our team ranked fifth out of the 11 participating teams after the Phase I evaluation.
@@ 2.2 Diarizen System
The Diarizen  [^4] https://github.com/BUTSpeechFIT/Diarizen/tree/main [12, 13, 14] system is the state-of-the-art SD system built on end-to-end neural diarization with vector clustering (EEND-VC) [15, 16, 17]. It is a hybrid speaker diarization pipeline. At the micro-level, it utilizes a powerful local end-to-end neural module, composed of a pre-trained WavLM [18] feature extractor, a Conformer [19] block, and a linear classifier, to process raw audio in short, overlapping chunks of typically $80$ to $16$ second windows. This acts as a neural frontend, which generates frame-level speech probabilities and extracts highly discriminative speaker embeddings while handling overlapping speech.
@@ 5 Conclusions and Future Directions
Although our best system ranked fifth among 11 unique teams on the leaderboard  [^8] Codabench DISPLACE-M Challenge Results of the DISPLACE-M Challenge under the Diarizen Evaluation Phase 1, we identify several limitations in our study and outline multiple promising directions for future work.